Ida Sidha Karya Company is a family-owned company located on the island of Bali in Indonesia. The company produces a handcrafted Balinese musical instrument called a gamelan that is similar to a xylophone. The gamelans are sold for $920. Selected data for the company’s operations last year follow:

Units in beginning inventory 0
Units produced 275
Units sold 260
Units in ending inventory 15
Variable costs per unit:
Direct materials $ 110
Direct labor $ 320
Variable manufacturing overhead $ 40
Variable selling and administrative $ 15
Fixed costs:
Fixed manufacturing overhead $ 77,000
Fixed selling and administrative $ 33,000

The absorption costing income statement prepared by the company’s accountant for last year appears below:

Sales $ 239,200
Cost of goods sold 195,000
Gross margin 44,200
Selling and administrative expense 36,900
Net operating income $ 7,300

Required:

1. Under absorption costing, how much fixed manufacturing overhead cost is included in the company's inventory at the end of last year?

2. Prepare an income statement for last year using variable costing.
Answer #1
1
Fixed manufacturing overhead cost included in inventory 4200 =77000/275*15
2
Variable costing Income Statement
Sales 239200
Variable expenses:
Variable cost of goods sold 122200 =260*(110+320+40)
Variable selling and administrative expenses 3900 =260*15
126100
Contribution margin 113100
Fixed expenses:
Fixed manufacturing overhead 77000
Fixed selling and administrative expenses 33000
110000
Net operating income 3100